IIT JEE Entrance Exam

230 JOULE WERE SPENT IN LIFTING A 10 KG WEIGHT TO A HEIGHT 2 m .CALCULATE THE ACCELERATION WITH WHICH IT WAS RAISED .TAKE g=10m/s2

omkar ajnadkar
14 Years agoGrade

2 Answers

nishad bagwe
14 Years ago

by conservation of energy,

the energy spent in doing the work=change in ke + change in pe

therefore, 230=10 x 2 x 2  + 0.5 x 10 x v2

230=40+5v2

v2=190/5

=38m/s

now,v2=2as

a=38/4

=9.5 m /s2

Sahil
8 Years ago
According to mechanical energy conservation :- 230 = 1/2*10*v^2 + 10 * 10 * 2230-200= 5v^2v^2 = 2as6/2*2=a= 1.5m/s^2